Skip to end of metadata
Go to start of metadata

Day 2 of IOP.  Weather fine and sunny with just widely scattered cirrus, almost no wind.  

Four soundings launch today (0Z, 6Z, 12Z & 18Z).  The office trailer for the sounding site was supposed to arrive today however was delayed again - this time apparently due to a truck breakdown.  They are now saying it will arrive tomorrow morning. In the meantime, we have no surface met or DM computer at the sounding site, and so no sounding data is being transferred. 

The profiler site was shutdown for several hours while the power company installed an extra line to enable DLR to operate their radiometer at the site.  We also installed an air conditioning unit in the shed, since the temperatures inside have been over 40C many days and occasionally reaching over 45C.

 

 

  • No labels